# MultiNash-PF: A Particle Filtering Approach for Computing Multiple Local Generalized Nash Equilibria in Trajectory Games

### v2

@@ V-1 Comparison with Baseline
In order to compare the effectiveness of our method, we consider a baseline method. In the baseline method, we directly initialize the IPOPT solver with random perturbations of reference trajectories and compute the resulting optimal solutions. For our method, we fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section IV. After that, we extract all the different equilibria and provide them as a warm starting point for IPOPT to obtain the final solutions. The compute time of our method is a combination of time used in implicit particle filtering and time used by IPOPT to obtain all the different equilibria given warm starting from filtering.
We run the IPOPT solver until we recover all 6 different equilibria. We repeat this experiment for 100 different Monte Carlo runs. We observe that it requires, on average, $18.64\pm 7.57$ number of IPOPT runs to obtain all six different equilibria using the baseline approach of random warm-starting. However, with our method, the number of IPOPT runs required is always six because we always obtain all six modes of equilibria from implicit particle filtering. A histogram of the number of IPOPT runs required for obtaining all six equilibria is plotted in Fig. 5(b).
Furthermore, the solve time to recover all 6 different Nash equilibria for the baseline is $6.71\pm 2.83s$, while for MultiNash-PF, the solve time is $3.36\pm 0.59s$. Our method results in a solve-time reduction of $50\%$ compared to the baseline while giving a much lower variance on the solve time. A histogram of the solve times of the baseline compared to our method is plotted in Fig. 5(a). As explained earlier, our method’s solve time is a combination of the time used in implicit particle filtering and the time used by IPOPT. We observe that the time required for implicit particle filtering over 100 Monte Carlo runs is $0.37\pm 0.072s$ while the time used by IPOPT to obtain the final solutions is $2.99\pm 0.57s$. Therefore, in our MultiNash-PF, the time used by implicit particle filtering is $11\%$ of the total solve time, which shows that filtering is very efficient in exploring the solution space of local Nash equilibria and contributes really less to the overall computation time of MultiNash-PF.
@@ VI Hardware Experiment
Finally, we conduct a hardware experiment where a human and a robot exchange their position as shown in Fig. 1. We perform these experiments using a TurtleBot 4 which is a ground robot with unicycle dynamics. We assume that the human also follows unicycle dynamics. The collision avoidance constraint between the TurtleBot and human is assumed to be the same as of (14). We use the Vicon motion capture system to measure the current position of both the robot and the human.
@@
Similar as shown in Fig. 2, two GNEs may exist which will correspond to two different modes of interactions. Namely, either the human and robot yield to their right or they yield to their left. For the robot to safely navigate around the human without collision, the robot needs to be aware of both the GNEs and needs to be able to determine which GNE the human chooses. Using MultiNashPF, the robot first uses an implicit particle filter to obtain the coarse estimates of the two different equilibria and then utilizes them as a warm start for IPOPT to obtain two modes of interaction. We assume that the interaction happens over $T=10s$. Then similar to [11], the robot uses a potential function based game theoretic planner in a model-predictive fashion with a planning horizon of 5s and $\Delta t=0.1s$ to plan its motion.
@@
The robot faces two choices of GNEs to follow recovered from the MultiNashPF. At any given time $t$ of planning, the robot computes the Fréchet distance [37] between the trajectory exhibited by the human so far and the human’s trajectory in both the computed GNEs. Let the distances of human trajectory from both the GNEs be denoted by $d_{1}$ and $d_{2}$, respectively. The robot keeps an uncertainty on the GNE that the human chooses until $|d_{1}-d_{2}|>d^{*}$ where $d^{*}$ is the specified threshold. For example, in Fig. 1 at $t=2$, values of $d_{1}$ and $d_{2}$ are almost identical which indicates that human has not chosen their mode of interaction. Whenever $|d_{1}-d_{2}|>d^{*}$ occurs, the robot becomes certain of the mode of interaction followed by the human. Then the robot follows GNE corresponding to $\arg\min_{i\in\{1,2\}}\{d_{i}\}$. For example, in Fig. 1, at $t=4$, the Fréchet distance of the human from its first GNE trajectory is very high while for the second GNE, the Fréchet distance is low indicating that the human has chosen the second GNE which corresponds to both the agents yielding to their right. Using our method, the robot autonomously identifies this mode, moves accordingly, and reaches its goal location while successfully avoiding collision with the human.
@@
We ran 10 different trials and each time we asked the human to either yield to their right or left randomly. Our algorithm was able to successfully compute both modes and identify the mode that the human was following. Using our algorithm, the robot was able to successfully avoid collision with the human in real-time and reach its goal location in all 10 runs.
@@ VII Conclusion
In this work, we introduced MultiNash-PF, a novel algorithm for efficiently computing multiple local GNEs in CPTGs. By leveraging potential game theory and implicit particle filtering, MultiNash-PF efficiently identifies coarse estimates of local minimizers, which are then refined using optimization solvers to obtain distinct interaction modes. Our numerical simulations demonstrate that MultiNash-PF significantly reduces computation time—by up to 50% compared to baseline methods—while effectively capturing the multimodal nature of multi-agent interactions. Furthermore, our real-world human-robot interaction experiments highlight the algorithm’s ability to reason about multiple interaction modes and resolve conflicts in real-time. These results underscore the potential of MultiNash-PF to enhance interactive motion planning, enabling more adaptive and socially aware autonomous systems.

### v3

@@ V-1 Comparison with Baseline
In order to compare the effectiveness of our method, we consider a baseline method. For the baseline, we directly initialize the IPOPT solver with random perturbations of reference trajectories and compute the resulting optimal solutions. The computation time of our method is a combination of the time used in implicit particle filtering and the time used by IPOPT to obtain all the different equilibria given warm starting from filtering.
For the baseline, we randomly initialize the IPOPT solver until we recover all 6 different equilibria. We repeat this experiment for 100 different Monte Carlo runs. We observe that it requires, on average, $18.64\pm 7.57$ random initializations of IPOPT to obtain all six different equilibria using the baseline approach. However, with our method, the number of IPOPT runs required is always six because we always obtain all six modes of equilibria from implicit particle filtering. A histogram of the number of IPOPT runs required for obtaining all six equilibria is plotted in Fig. 5(a).
Furthermore, the solve time to recover all 6 different Nash equilibria for the baseline is $6.71\pm 2.83s$, while for MultiNash-PF, the solve time is $3.36\pm 0.59s$. Our method results in a solve-time reduction of $50\%$ compared to the baseline while giving a much lower variance on the solve time. A histogram of the solve times of the baseline compared to our method is plotted in Fig. 5(b). As explained earlier, our method’s solve time is a combination of the time used in implicit particle filtering and the time used by IPOPT. We observe that the time required for implicit particle filtering over 100 Monte Carlo runs is $0.37\pm 0.072s$ while the time used by IPOPT to obtain the final solutions is $2.99\pm 0.57s$. Therefore, in our MultiNash-PF, the time used by implicit particle filtering is $11\%$ of the total solve time, which shows that filtering is very efficient in exploring the solution space of local Nash equilibria.
@@ VI Hardware Experiment
Finally, we conduct a hardware experiment where a human and a robot exchange their positions as shown in Fig. 1. We perform these experiments using a TurtleBot 4, which is a ground robot with unicycle dynamics. We assume that the human also follows unicycle dynamics. The collision avoidance constraint between the TurtleBot and the human is assumed to be the same as that of (14). We use the Vicon motion capture system to measure the current position of both the robot and the human.
@@
Similar as shown in Fig. 4, two GNEs may exist, which will correspond to two different modes of interactions. Namely, yield-right equilibrium when the human and robot yield to their right and yield-left equilibrium when they yield to their left. For the robot to safely navigate around the human without collision, the robot needs to be aware of both the GNEs and needs to be able to determine which GNE the human chooses. Using MultiNash-PF, the robot first uses an implicit particle filter to precompute the coarse estimates of the two different equilibria and then utilizes them as a warm start for IPOPT to obtain two modes of interaction. We assume that the interaction happens over $T=10s$. Then, similar to [11], the robot uses a potential function based game-theoretic planner in a model-predictive fashion with a planning horizon of 5s and $\Delta t=0.1s$ to plan its motion.
@@
The robot faces two precomputed choices of GNEs to follow recovered from the MultiNash-PF. At any given time $t$ of planning, the robot computes the Fréchet distance [38] between the trajectory exhibited by the human so far and the human’s trajectory in both the computed GNEs. Let the distances of the human trajectory from both the GNEs be denoted by $d_{1}$ and $d_{2}$, respectively. The robot keeps an uncertainty on the GNE that the human chooses until $|d_{1}-d_{2}|>d^{*}$ where $d^{*}$ is a pre-specified threshold. For example, in Fig. 1, at $t=2$, the values of $d_{1}$ and $d_{2}$ are almost identical, which indicates that the human has not chosen their mode of interaction. Whenever $|d_{1}-d_{2}|>d^{*}$ occurs, the robot becomes certain of the mode of interaction followed by the human. Then, the robot follows the GNE corresponding to $\arg\min_{i\in\{1,2\}}\{d_{i}\}$. For example, in Fig. 1, at $t=4$, the Fréchet distance of the human from its yield-left equilibrium trajectory is very high, while for the yield-left equilibrium, the Fréchet distance is low, indicating that the human has chosen the yield-right equilibrium. Using our method, the robot autonomously identifies this mode, moves accordingly, and reaches its goal location while successfully avoiding collision with the human.
@@
We ran 10 different trials, and each time, we asked the human to either yield to their right or left randomly. Our algorithm was able to successfully compute both modes and identify the mode that the human was following. Using our algorithm, the robot was able to successfully avoid collision with the human in real-time and reach its goal location in all 10 runs.
@@ VII Conclusion
In this work, we introduced MultiNash-PF, a novel algorithm for efficiently computing multiple interaction modes. By leveraging potential game theory and implicit particle filtering, MultiNash-PF efficiently identifies coarse estimates of interaction modes, which are then refined using optimization solvers to obtain distinct interaction modes. Our numerical simulations demonstrate that MultiNash-PF significantly reduces computation time by up to 50% compared to baseline methods, while effectively capturing the multimodal nature of multi-agent interactions. Furthermore, our real-world human-robot interaction experiments highlight the algorithm’s ability to reason about multiple interaction modes and resolve conflicts in real-time.